OpenStudy (anonymous):

(4y/y+3)+(y/y-3) please help with the fractions!

OpenStudy (anonymous):

I think its 5y^2-9y but not sure "/

OpenStudy (anonymous):

the problem is, =4y/(y+3)+y/(y-3) =[4y(y-3)+y(y+3)]/(y+3)(y-3) =[4y^2-12y+y^2+3y]/(y^2-3^2) =[5y^2-9y]/[y^2-9]

OpenStudy (anonymous):

no hes lying dont trust him

OpenStudy (anonymous):

dont trust you. either

OpenStudy (anonymous):

I saw this big yo mama fight and he totally broke up the whole thing

OpenStudy (anonymous):

THE YO MAMA PPL IS INSULTING ME so i had to break up the whole thing doesn't mean u can't trust me. anyway i never been this famous before just because the YO MAMA thing.

OpenStudy (anonymous):

you read that so plz be qiuet about it @sman493